HIGH-RESOLUTION LASER SPECTROSCOPY OF A^{3}Π_{1u} ← X^{1}Σ_{g}^{+} SYSTEM OF I_{2}

\underline{N. NISHIMIYA}, T. YUKIYA, O. OHSHIMA, AND M. SUZUKI, Dept. of Electronic Engineering, Tokyo Institute of Polytechnics, Iiyama 1583, Atsugi City, 243-0297 Kanagawa, Japan.

The molecular constants for the A-state of I_{2} calculated by the NDE method have been reported by D.R.T.Appadoo et al.^{1} Yukiya et al. measured the Q-branch lines up to v′ = 45 using a titanium sapphire ring laser, where the hyperfine components were partially resolved at the higher vibrational bands than those of v′ = 35, and determined the spectroscopic constants.^{2} In this work we report the line splittings of the P and R-branches caused by the nuclear quadrupole coupling effect. The 23 bands of (v′ = 15 − 19) ← (v′′ = 3), (v′ = 18 − 30) ← (v′′ = 4), and (v′ = 25 − 29) ← (v′′ = 5) were newly assigned. The unperturbed line positions were obtained using the high J approximation. The spectroscopic constants of T_{v}, B_{vf}, D_{v}, H_{v}, and q_{v} of the A-state determined by the aid of the Dunham parameters of the X-state^{3} will be presented.

1. D.R.T.Appadoo, R.J.LeRoy, P.F.Bernath, S.Gerstenkorn, P.Luc, J.Vergès, J.Sinzelle, J.Chevillard, and Y.D’Aignaux, J.Chem.Phys., 104, 903 (1996).

2. T.Yukiya, N.Nishimiya, and M.Suzuki, J.Mol.Spectrosc., 182, 271 (1996).

3. S.Gerstenkorn and P.Luc, J.Phys., 46, 867 (1985).